OpenStudy (anonymous):

Find the x- and y-intercepts for the following equation. −2x + 3y = 15 x-intercept (x, y) = y-intercept (x, y) = ,

OpenStudy (anonymous):

you can find the x intercept by subbing in 0 for y and you can find the y intercept by subbing in 0 for x. x intercept : -2x + 3y = 15 -2x + 3(0) = 15 -2x = 15 (divide by -2) x = -15/2 x intercept = (-15/2,0) y intercept : -2x + 3y = 15 -2(0) + 3y = 15 3y = 15 (divide by 3) y = 5 y intercept = (0,5)

OpenStudy (anonymous):

how do i graph -15/2 , 0

OpenStudy (anonymous):

you can turn it into a mixed number....-15/2 = - 7 1/2...that will make it easier to graph

OpenStudy (anonymous):

so your x intercept is (-7 1/2, 0) if that makes it easier to graph

OpenStudy (anonymous):

Find the x- and y-intercepts for the following equation. y = 2x + 4 x-intercept (x, y) = (, ) y-intercept (x, y) = (, )

OpenStudy (anonymous):